# Robust composite Molmer-Sorensen gate

### v1

@@ V Discussion and conclusion
In this work we presented the combined technique of amplitude modulation and composite gates to produce a Molmer-Sorensen gate which is robust to timing errors, rotational errors and detuning errors. To our knowledge, the sine-cosine amplitude modulation is new and individually it performs quite well against both timing and detuning errors. The permuted sequence of Eq. (18) that we propose negates the effects of errors propagation from the two-qubit part of the sequence to the single-qubit gates in it and maintains the robustness of the amplitude modulation versus gate timing errors. The $\sin(3t/2)^{2}$ modulated sequence outperforms the two-tone excitation technique versus all types of errors except for an error in the detuning, although its performance there is quite similar. It also surpassed the sine-cosine modulation, which performs better individually. Based on that we can conclude that the best performing modulation individually is not necessarily the best performing modulation in a sequence. Furthermore, the $\sin(t/4)^{2}$ modulation’s trajectory is very similar to the two-tone’s trajectory, yet their performances, both individually and in a sequence, differ strongly. Thus a clear indication for the performance of a MS gate, based only on its phase-space trajectory can not be derived. There is a clear trade-off between time, power and robustness which is an expected effect, also observed in single qubit gates [29]. In general, the power costs of a sequence with modulated Rabi frequency are a bit below 2.5 times the cost of a single MS gate with the same modulation. Different modulations allow for a variety of gate times ranging from the minimal gate time for a single MS gate to a few multiples of it. However, choosing a proper signal for modulation in a composite gate has to be in accordance with the power capabilities of the implementing platform and its individual tolerances to errors of different nature.
@@ Acknowledgements.
This research is supported by the Bulgarian national plan for recovery and resilience, Contract No. BG-RRP-2.004-0008-C01 (SUMMIT), Project No. 3.1.4 and by the European Union’s Horizon Europe research and innovation program under Grant Agreement No. 101046968 (BRISQ).
Appendix A Analytical expressions for $\alpha$ and $\theta$
@@
.$\Bigg{.}-2\epsilon\sin\left(t/2\right)\bigg{(}1-\cos(t\epsilon)+32\epsilon^{4}-12\epsilon^{2}\bigg{)}+t\epsilon-\sin(t\epsilon)\Bigg{]}.$
@@
.$\Bigg{.}\left.-24\left(6\epsilon^{4}-23\epsilon^{2}+8\right)\sin(t)+24\left(2\epsilon^{4}-7\epsilon^{2}+2\right)\sin(2t)\Bigg{]}+96\epsilon\left(\epsilon^{2}+2\right)\sin(t)\left(2\left(\epsilon^{2}-1\right)\cos(t)-\epsilon^{2}+4\right)\cos(t\epsilon)\right.$ (20b)
@@
In a similar fashion $\alpha$ and $\theta$ can be derived for the time intervals of the sequences in Section IV.

### v3

@@ V Discussion and conclusion
In this work we presented the combined technique of amplitude modulation and composite gates to produce a Molmer-Sorensen gate which is robust to timing errors, rotational errors and detuning errors and can also withstand certain levels of a.c. Stark shifts and drifting detuning. To our knowledge, the sine-cosine amplitude modulation is new and individually it performs quite well against both timing and detuning errors. The permuted sequence of Eq. (20) that we propose negates the effects of errors propagation from the two-qubit part of the sequence to the single-qubit gates in it and maintains the robustness of the amplitude modulation versus gate timing errors. The $\sin(3t/2)^{2}$ modulated sequence outperforms the two-tone excitation technique versus all types of errors except for an error in the detuning, although its performance there is quite similar. It also surpassed the sine-cosine modulation, which performs better individually. This is due to errors in the displacement operator. We came to this conclusion by comparing the behaviour of the sequences under the assumption of perfect displacement operators which are errorless then the behaviour is quite similar to the single gate case where the $\sin^{2}(t/2)\cos(t)$ signal is the most robust one. Individually, the $\sin^{2}(t/2)\cos(t)$ modulation outperforms the rest of the signals if its trajectory is allowed to complete a full traversing of the phase space and end back in the origin. If we interrupt it during the interaction, in order to speed-up the procedure the over-all error increases. This is due to the $\cos(t)$ part of the modulation, since it is responsible for the wings of the pulse, which generate the small arcs in the beginning and by the end of the evolution in phase space(see Fig. 6 b)). During the evolution the curve acquires a phase which at specific intervals rephases the trajectory in a similar way to dynamical decoupling. For the $\sin^{2}(t/2)\cos(t)$ modulation this effect is most profound for the wings of the pulse, therefore disrupting the trajectory loses the strength of the technique. Based on that we can conclude that the best performing modulation individually is not necessarily the best performing modulation in a sequence. Furthermore, the $\sin(t/4)^{2}$ modulation’s trajectory is very similar to the two-tone’s trajectory, yet their performances, both individually and in a sequence, differ strongly. Thus a clear indication for the performance of a MS gate, based only on its phase-space trajectory can not be derived. There is a clear trade-off between time, power and robustness which is an expected effect, also observed in single qubit gates [32]. In general, the power costs of a sequence with modulated Rabi frequency are a bit below 2.5 times the cost of a single MS gate with the same modulation. Different modulations allow for a variety of gate times ranging from the minimal gate time for a single MS gate to a few multiples of it. However, choosing a proper signal for modulation in a composite gate has to be in accordance with the power capabilities of the implementing platform and its individual tolerances to errors of different nature.
@@ Acknowledgements.
This research is supported by the Bulgarian national plan for recovery and resilience, Contract No. BG-RRP-2.004-0008-C01 (SUMMIT), Project No. 3.1.4 and by the European Union’s Horizon Europe research and innovation program under Grant Agreement No. 101046968 (BRISQ). We are greatful to the following projects VU-F-205/2006,DO-02-136/2008,DO-02-167/2008,DO 02-90/2008 for granting access to the PHYSON cluster.
Appendix A Amplitude Modulation
Here we introduce the technique of amplitude modulation for two-qubit gates. Ultimately our goal is to combine it with composite pulses keep their benefits while diminishing their flaws. We will employ amplitude modulation for its ability to mitigate timing and detuning errors and we will also join together multiple gates, generating a sequence, mimicking composite pulses techniques, to eliminate rotational errors. In order to find a proper time-dependent Rabi frequency which will be able to produce the MS gate, there are a couple of assumptions we can make based on Eqs. (3). First, we demand that the $M_{1}(t)$ term evaluates to 0 at the gate time which will ensure a closed trajectory in the phase space. Second, due to the form of the $M_{2}(t)$ term, the Rabi frequency must have a good auto-correlation in order to achieve a strong excitation with minimal amount of power. Both of these assumptions have to be satisfied simultaneously. A natural candidate functions that satisfy these requirements are the trigonometric functions. Using a sine function provides a smooth-start [31], which minimizes the coupling to the carrier and also has proven to have a low power requirement [20]. In order to find a larger family of functions, we can combine it with a cosine of a certain degree, as long as the smooth start is ensured. In its most general form the coupling of Eq. (1) then reads
$g(t,m,l,n,p)=A\sin(mt)^{l}\cos(nt)^{p}.$ (22)
Such amplitude modulation provides analytical expressions for $\alpha(t)$ and $\theta(t),$ for all different sets of parameters $m,l,n,p.$ They are too cumbersome to be presented here so instead we demonstrate the dynamics of a few of these modulations shown Fig. 6. We present their analytical form used for this illustration in the Appendix B. These modulations generate a variety of trajectories, illustrated in Fig. 6(b), some quite similar and others quite different. The relation between the phase space trajectory and the robustness versus specific parameter is not fully understood. There are empirical observations which state that on average the trajectory must be as close to the center as possible. A strict mathematical derivation describing the relation between robustness and trajectory is yet to be derived. We note that there are attempts to achieve this with an interesting algebraic approach  [20]. We now turn our attention to the selected modulations and point out that since the parameter space is quite rich the following discussion is not aimed at drawing the borders of the best performing parameter region, but rather to draw quantitative conclusions based on the given examples. For example, the sine-squared shape modulation is in general quite good for gate timing errors, as it transforms the point of entanglement into a plateau. Changing the frequency of the sine also seems to impact the infidelity, as can be seen from Fig. 6(c). When we look at the performance against detuning errors, shown in Fig. 6(d), we see that the $\sin(t/2)^{2}\cos(t)$ outperforms the other sine shapes. Note the asymmetry in Fig. 6(d), which is due to a resonance. In this example of amplitude modulation a larger detuning is required so that a closed trajectory can be formed and consequently a stronger coupling is needed. Thus a clear trade-off emerges between gate-time, power and robustness to multiple parameters, that one has to keep in mind, when choosing a specific modulation signal with respect to the experimental platform at hand.
Refer to caption (a) Refer to caption (b)
Refer to caption (c) Refer to caption (d)
In our simulations we use only two-tone excitation optimized for detuning errors, and for that reason it gains infidelity versus gate timing errors. We also point, however, that it can be optimized for gate timing errors as in Ref. [19]. We see that for detuning errors the sine-cosine modulation can underperform or outperform the multi-tone excitation depending on the sign of the detuning error. In any case it is more advantageous to use such amplitude modulation since it also provides gate timing robustness. We will use it in a composite sequence in order to generate a gate that is robust against any type of error. As well as the other two signals, namely the $\sin(t/4)^{2}$ and the $\sin(3t/2)^{2}$ modulations. Thus we attempt to draw any conclusion about their performance based on their phase-space trajectory, as the former provides quite a different trajectory than the multi-tone excitation, while the later has its trajectory quite similar. In terms of fidelity the former performance is somewhat in the middle of all other shapes, while the later has the worst performing infidelities for gate timing errors. It is worthwhile to point out that these two modulations behave even worse than the standard gate versus detuning errors, and it is crucial to clarify that this behavior is due to the phase-space trajectory, which we can also change by the detuning, rather than the modulation itself.
Appendix B Analytical expressions for $\alpha$ and $\theta$
@@
$\Bigg{.}-2\epsilon\sin\left(t/2\right)\bigg{(}1-\cos(t\epsilon)+32\epsilon^{4}-12\epsilon^{2}\bigg{)}+t\epsilon-\sin(t\epsilon)\Bigg{]}.$
@@
$\Bigg{.}\left.-24\left(6\epsilon^{4}-23\epsilon^{2}+8\right)\sin(t)+24\left(2\epsilon^{4}-7\epsilon^{2}+2\right)\sin(2t)\Bigg{]}+96\epsilon\left(\epsilon^{2}+2\right)\sin(t)\left(2\left(\epsilon^{2}-1\right)\cos(t)-\epsilon^{2}+4\right)\cos(t\epsilon)\right$. (24b)
@@
In a similar fashion $\alpha$ and $\theta$ can be derived for the time intervals of the sequences in Section III.